i need to solve this equation -2x^2+2x+4=-8 how do i do it
Answer by stanbon(75887)

solve this equation
-2x^2+2x+4=-8
----
Divide thru by -2 to get:
x^2 -x -2 = 4
---
x^2 -x -6 = 0
Factor:
(x-3)(x+2) = 0
x = 3 or x = -2
=====================
cheers,
Stan H.